Shift report 20020104 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Fri Jan 04 2002 - 08:05:14 EST

  • Next message: Flemming Videbaek: "Fw: new paper from Phobos"

    ------------------------------------------------------------
    Date: Fri Jan  4 08:05:14 2002, Shift: 00:00-08:00
    Supervisor: hiro, with: 
    Log book#11 pages 145 - 145
    
    Shift Summary:
    ------------------------------------------------------------
    1.
    00:30  Turn on all magnets
    01:15  New beam, but it is noticed that D2 and D5 magnet 
           have been tripped.  Call the magnet watch.
    
    01:45  Magnet watch fixed D2 and D5.
    
    01:54  New Run 6270
    
           Check Trig 3 to local track efficiency requested by FV
           Out of 30 Trig 3 events:
    	   TPM1 had 23 events with track:23/30=77%
    	   TPM2 had 21 events with track:21/30=70%
    
    03:57  New Run 6271
           It is noticed that Trig6 is not working quite right.
    
    07:20  MCR called to dump beam.  End the Run
    
    
    2.  Run6270, 6271
    3.  NONE detected.
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Fri Jan 04 2002 - 08:06:40 EST